OpenStudy (anonymous):

can someone show me how to find the min value for a function and show me how to determine the function's minimum y-value.

OpenStudy (welshfella):

by using Calculus?

OpenStudy (anonymous):

im in trigonometry at the moment

OpenStudy (anonymous):

have you studied derivatives?

OpenStudy (welshfella):

well there are known minimums of trig functions eg for sine and cosine it is -1 so if you need to find the minimum value of 3 cos x it is -3

OpenStudy (anonymous):

so if i had a function g(x) = 2 sin(2x - π) + 4 2sinx=-2

OpenStudy (welshfella):

minimum of 2 sinx is -2, yes

OpenStudy (anonymous):

ok great thnx

OpenStudy (anonymous):

how do you find the min of a regular function

OpenStudy (welshfella):

what do you mean by a regular function ? an algebraic function like say x^3 + x^2 - 5x - 11?

OpenStudy (anonymous):

yeah

OpenStudy (welshfella):

you need calculus to do that

OpenStudy (anonymous):

f(x) = 2x^3+ 16x + 12

OpenStudy (anonymous):

what abut for this one

OpenStudy (welshfella):

you need to find the derivative f'(x)

OpenStudy (anonymous):

can you show me how

OpenStudy (welshfella):

its too much to show you in one post if you haven't done differentiation yet

OpenStudy (anonymous):

oh ok but thnx for the help though one last question about the trig func. though how do you find the min y-value

OpenStudy (welshfella):

example y = 2 cos x + 3 minimum y value = 2*-1 + 3 = 1

OpenStudy (anonymous):

thank you sooo much your were the biggest help :3